Find the distance between the given numbers
a.) 2 and 17

Answer by eperette(173)

Concept you should know:
(1) distance is always a positive number
(2) distance = abs(a-b), (i am using abs for absolut value)
Given 2 and 17
Answer:
distance = abs( 2- 17)
distance = abs(-15)
distance = 15
.....Note that you should have received same answer by doing the following:
distance = abs (17-2)
distance = abs (15)
distance = 15